OpenStudy (rock_mit182):

solve the following problem by using the Quadratic Formula 5w^2 - 3

OpenStudy (rock_mit182):

i don't know how to and the second term, in order to apply the cuadratic formula

OpenStudy (rock_mit182):

how to find *

OpenStudy (jdoe0001):

\(\bf 5w^2 - 3\implies 5w^2 +0w- 3\\ \textit{quadratic formula}\\ {\color{blue}{ 5}}w^2{\color{red}{ +0}}w{\color{green}{ -3}} \qquad \qquad w= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (rock_mit182):

what about : \[(\sqrt{5}w + \sqrt{3}) (\sqrt{5}w - \sqrt{3})\] can I do this ?

OpenStudy (jdoe0001):

you could factor it, yes, however -> "olve the following problem by using the \(\bf Quadratic\ Formula\)"

OpenStudy (rock_mit182):

but that would be an imaginary solution

OpenStudy (jdoe0001):

yes.. and factoring will give the same anyhow...but you're asked to use the quadratic formula

OpenStudy (rock_mit182):

ok so that woulde all in this case: 5w^2-3 = 0

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

"solving" a quadratic equation, means just finding the x-intercepts and thus setting y = 0

OpenStudy (rock_mit182):

ok Im going to do the whole process and then call yo just to check it out

OpenStudy (jdoe0001):

ok

OpenStudy (rock_mit182):

i was thinking if instead just... w^2 = 3 \[w =\pm \sqrt{3}\]

OpenStudy (rock_mit182):

i have to get the same result using the quadratic ?

OpenStudy (jdoe0001):

yeap, simplifying or factoring, or using the quadratic, will yield the same result

OpenStudy (rock_mit182):

i meant : \[w = \pm \sqrt{\frac{ 3 }{ 5 }}\]

OpenStudy (rock_mit182):

ok

OpenStudy (rock_mit182):

\[x = \frac{ -(0) \pm \sqrt{0^{2}-4(5)(-3)} }{ 2(5) }\]

OpenStudy (rock_mit182):

\[x =\frac{ \pm \sqrt{-4(-3)(5)} }{ 10 }\]

OpenStudy (rock_mit182):

\[x =\frac{ \pm \sqrt{60} }{ 10 }\]

OpenStudy (rock_mit182):